### v1

# Long-term Frame-Event Visual Tracking: Benchmark Dataset and Baseline

Abstract
Current event-/frame-event based trackers undergo evaluation on short-term tracking datasets, however, the tracking of real-world scenarios involves long-term tracking, and the performance of existing tracking algorithms in these scenarios remains unclear. In this paper, we first propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and has become the largest frame-event tracking dataset to date. We re-train and evaluate 15 baseline trackers on our dataset for future works to compare. More importantly, we find that the RGB frames and event streams are naturally incomplete due to the influence of challenging factors and spatially sparse event flow. In response to this, we propose a novel associative memory Transformer network as a unified backbone by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on both FELT and RGB-T tracking dataset LasHeR fully validated the effectiveness of our model. The dataset and source code can be found at https://github.com/Event-AHU/FELT_SOT_Benchmark.
Keywords:
RGB-DVS tracking Benchmark Dataset Transformer Modern Hopfield Network
1 Introduction
Visual object tracking aims to locate the specified object using an adaptively adjusted bounding box. This task has been widely employed in practical applications, such as automatic piloting, drone aerial photography, and intelligent security monitoring. Usually, the visual trackers are developed and deployed based on RGB cameras which record the scenario using video frames. As we know, RGB cameras are easily influenced by challenging factors, such as low illumination, fast motion, heavy occlusion, etc. Therefore, tracking by using RGB cameras is still a hard task even nowadays.
To address the aforementioned issues, the researchers begin to exploit new sensors to replace or assist the RGB cameras for more robust tracking. Among them, bio-inspired event cameras (DVS346, CeleX-V, Prophesee, etc) have drawn more and more attention in recent years due to their unique imaging principles. Different from RGB cameras which record all the pixel intensities in the scene in a synchronous way and output video frames at a fixed frame rate (e.g., 30 frames per second), the event cameras emit a spike or event point for each pixel only when the variation of light intensity exceeds a given threshold. The event pixel values (-1 or 1) are independently and asynchronously recorded, and -1 or 1 (corresponding to an OFF or ON event) denotes that the light intensity is decreased or increased. As noted in [9], event cameras also perform better than the RGB cameras on the high dynamic range, low energy consumption, and high temporal resolution. Thus, it can be used in low-illumination and fast-motion tracking scenarios. Even with the above advantages, relying solely on event cameras still performs poorly in some scenes, such as static or motionless objects, because only the signals of motion can be sensed, and stationary objects cannot be captured. Event cameras also fail to capture the color information or very fine-grained texture information. Fortunately, this is where the RGB cameras excel. Therefore, the usage of both RGB and event cameras for tracking has great landing value and potential applications.
Based on this idea, there are already several datasets proposed for event-based tracking, including FE108 [33], VisEvent [23], COESOT [22], and EventVOT [24]. Specifically, the FE108 [33] is collected using a gray-scale DVS event camera in the indoor scenarios and EventVOT [24] is a pure event-based tracking dataset. VisEvent [23] and COESOT [22] are two large-scale tracking datasets, however, most of their samples are short-term. Conducting long-term frame-event single object tracking has the following advantages and benefits: 1) Long-term frame-event tracking is closer alignment with real-world applications; 2) Better handling frequent target changes, noise, and uncertainty; 3) Long-term frame-event tracking can also support the long-term monitoring of energy consumption. Based on the aforementioned observations and reflections, in this work, we propose the first benchmark dataset for Frame-Event Long-Term single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and has become the largest frame-event tracking dataset to date, as shown in Table 1. 14 attributes and 45 classes of target objects are involved in this dataset. In addition, we re-train and evaluate 15 baseline trackers on our dataset to achieve a fair comparison. We believe the dataset and benchmark results build a good platform for the development of the long-term frame-event single object tracking.
In this work, we also propose a novel frame-event tracking framework, termed AMTTrack. The key insight of our tracker is that the long-term tracking videos present the characteristics of incomplete information. For example, the RGB frames can be influenced by low illumination, motion blur, occlusion, etc, meanwhile, the event data is spatially sparse or with no target signal at all. This uncertain, incomplete, and not always complementary frame-event signal makes conventional fusion methods perform poorly. In this work, we build a unified backbone based on the modern Hopfield layer [19] which makes the ViT backbone have the ability of associative memory. Thus, the spatial-temporal information can be fully mined and learned. We integrate this strategy into the CEUTrack [22] and obtain higher performance on long-term frame-event tracking and RGB-T tracking (LasHeR dataset is evaluated). The overall framework is illustrated in Fig. 3.
To sum up, we draw the following contributions of this paper as the following three aspects:
  • We propose the first frame-event long-term single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and will be a better platform for the development of long-term frame-event-based tracking.
  • We propose a novel frame-event visual tracking framework based on the associative memory Transformer network, termed AMTTrack. It is the first work to introduce the modern Hopfield layer [19] to boost the memory capability of Transformer for the long-term frame-event tracking.
  • To assess existing event-based trackers and provide extensive baselines for future works to compare on the FELT SOT dataset, we re-train and evaluate 15 representative trackers under different modalities, including RGB, Event and RGB+Event.
2 Related Work
In this section, we review the prominent visual trackers based on Event/RGB-Event cameras, Event-based datasets, and Modern Hopfield Networks. More related works can be found on the paper list 111github.com/wangxiao5791509/Single_Object_Tracking_Paper_List.
2.1 Event Camera based Tracking
As a newly arising research topic in computer vision, Event-based tracking [33] quickly draws more and more attention. Wang et al. [23] propose the Cross-Attention Transformer (CMT) to enable feature interaction between RGB and event data for tracking. AFNet [32] proposed by Zhang et al. introduces a multi-modal alignment and fusion module, which adaptively merges complementary information from high frame rate event streams and RGB frames while attenuating noise. Wang et al. propose HDETrack [24] which utilizes hierarchical knowledge distillation to transfer multi-modal/view knowledge to the single-modal event stream, enabling high-speed and low-latency visual tracking. Tang et al. propose CEUTrack [22] which employs a unified Transformer backbone to model the relationship between RGB frames and event voxels, thereby facilitating effective feature interaction. Zhu et al. [40] introduce a mask modeling strategy aimed at addressing the issue of cross-modal interaction between RGB frames and event streams. Fu et al. introduce DANet [8], crafting an event-based interference sensing tracker through the integration of Transformer modules and Siamese architecture. ViPT [36] is proposed by Zhu et al. which incorporates learnable modal-relevant prompts while retaining frozen pre-trained models, enhancing the adaptability of the models to diverse multi-modal tracking tasks. Zhu et al. [39] directly process event clouds and predict the motion-aware target likelihood. Zhang et al. propose a cross-domain attention fusion algorithm in STNet [31]. Different from these works, in this work, we propose a novel long-term frame-event tracker using the associative memory Transformer.
Table 1: Comparison of Event camera-based visual tracking datasets.
Datasets Year Videos Frames Class Attributes Resolution Aim Absent Frame Reality LT Public
VOT-DVS 2016 60 - - - $240\times 180$ Eval
TD-DVS 2016 77 - - - $240\times 180$ Eval
Ulster 2016 1 9,000 - - $240\times 180$ Eval
EED 2018 7 234 - - $240\times 180$ Eval
FE108 2021 $108$ 208,672 21 4 $346\times 260$ Train & Eval
VisEvent 2021 $820$ 371,127 - 17 $346\times 260$ Train & Eval
COESOT 2022 1354 478,721 90 17 $346\times 260$ Train & Eval
EventVOT 2023 1141 569,359 19 14 1280 $\times$ 720 Train & Eval
FELT (Ours) 2024 742 1,594,474 45 14 $346\times 260$ Train & Eval
2.2 Event-based Benchmark Dataset
Event stream-based visual tracking has garnered increasing attention in recent years, highlighting the growing importance of event-based benchmark datasets. To uphold dataset quality, manual labeling of target boxes is often a requisite for event-based benchmark datasets. Existing widely used event-based single object tracking datasets include VisEvent [23], FE108 [33], COESOT [22], and EventVOT [24].
FE108 [33] comprises 108 videos captured under consistent lighting conditions, totaling 208,672 frames for the entire dataset, averaging 1,932 frames per video. The dataset encompasses 21 distinct categories of target objects and introduces four challenging factors.
VisEvent [23] encompasses 820 video sequences, totaling 371,127 frames, with an average length of 453 frames per video. Six fundamental guidelines were adhered to during the dataset collection process to ensure high-quality, diverse, and large-scale datasets. The dataset incorporates 17 challenging factors, encompassing RGB mode challenges such as fast motion, low illumination, and exposure, as well as event mode challenges like the absence of target movement, background confusion, etc.
COESOT [22] is a large-scale benchmark dataset for color-event tracking. It comprises 1,354 video sequences encompassing 90 target object categories. The dataset spans a total of 478,721 frames, averaging 354 frames per video. The diverse video scenes include indoor, street, zoo, and various other environments. Noteworthy challenges within the dataset involve factors such as the rotation of target objects, complete occlusion, and movement of background objects.
EventVOT [24] stands as the pioneering high-resolution, large-scale, single-modal event dataset. Encompassing 1,141 video sequences, the dataset comprises a total of 569,359 frames, averaging 499 frames per video. Capturing a diverse range of scenes, including indoor, street, and stadium, the videos span both day and night time periods. The dataset prominently features 14 challenging factors, such as speed changes and variations in lighting conditions. Notably, all data undergo annotation by a professional labeling company to ensure the dataset’s overall high quality.
Compared with these datasets, our proposed FELT is the first long-term frame-event tracking dataset which is the largest one, as shown in Table 1.
2.3 Modern Hopfield Networks
Hopfield network [12] proposed by John Hopfield in 1982 is well known for its associative memories and it is treated as one of the earliest artificial neural models. This neural network can associate an input with its most similar pattern, i.e., retrieve a stored memory based on partial or incomplete input. Different from the classical binary Hopfield networks, recent works [13, 7] demonstrate that by introducing polynomial or exponential terms into the energy function, the storage capacity can be increased significantly. Modern Hopfield networks [19]  [^2] https://ml-jku.github.io/hopfield-layers/ proposed by Ramsauer et al. achieves continuous associative memory. After that, many tasks are boosted with the help of Modern Hopfield networks, such as remote sensing image generation [27], reaction template prediction [21], immune repertoire classification [25]. Inspired by the strong associative memory of modern Hopfield networks  [^3] https://github.com/Event-AHU/Awesome_Modern_Hopfield_Networks, in this work, we propose the associative memory Transformer for long-term frame-event tracking by introducing them into the multi-head self-attention blocks.
3 FELT Single Object Tracking Dataset
3.1 Design Principle
We obey the following principles when collecting the FELT SOT dataset: 1). Long-term: Each video contains at least 1000 video frames and event streams. 2). Large-scale: To bridge the gap between data-hunger deep neural networks, we aim to build the largest frame-event tracking dataset. 3). Various Challenges: The collected videos need to reflect the key challenges in frame-event tracking. 4). Dual-Modality: To meet the demand for event-only and frame-event fusion-based tracking tasks, we need to build a dataset with both RGB and Event data. 5). Incomplete Information: Considering the particularity of multi-modal tracking, the dataset proposed in this paper needs to take into account the incomplete information of each modality. For example, the event data in spatial sparse and RGB data in the weakened scenarios.
Figure 1: (a). The number of videos corresponding to specific challenging factors; (b). Comparison between existing event-based tracking datasets and our proposed FELT dataset; (c). Visualization of the categories of target objects in our dataset.
3.2 Dataset Collection and Annotation
The DVS346 event camera is adopted for the collection of our FELT dataset, which outputs the RGB frames ($346\times 260$) and event streams simultaneously. The two modalities are aligned well based on hardware perspective on both spatial and temporal views. When recording our videos, we follow the design principles and challenging attributes defined in sub-section 3.1 and 3.3. For the bounding box annotation, we employ a professional data company to annotate and check for several rounds to ensure the quality.
3.3 Attributes
When collecting the FELT SOT dataset, we consider 14 attributes, as shown in Table 2, for the long-term frame-event tracking task. Specifically, the ST (Small Target) and illumination related attributes (OE (Over Exposure), IV (Illumination Variation), LI (Low Illumination)) are all considered. We also define the motion-based attributes, like the NMO (No MOtion) and FM (Fast Motion). The OV (Out of View) is considered for the long-term tracking task. We believe these defined challenging attributes will be a better platform for the evaluation of future trackers.
Table 2: 14 attributes defined in our proposed FELT SOT dataset.
Attributes Description
01. ST (Small Target) The scale of the target object is limited
02. OE (Over Exposure) The light intensity is very high
03. IV (Illumination Variation) The light intensity changes during tracking
04. NMO (No Motion) The target object is stationary
05. ARC (Aspect Ratio Change) The ratio of the bounding box aspect ratio varied significantly
06. BI (Background Influence) The target object is heavily influenced by the background
07. SV (Scale Variation) The width and height of the target object changed significantly
08. POC (Partial Occlusion) Part of the target object is occluded
09. OV (Out of View) The target object moves out of the view of the event camera
10. LI (Low Illumination) The videos are recorded in the dark environment
11. FOC (Fully Occlusion) The target object is fully occluded by other objects
12. DEF (Deformation) The shape of the target object changed during tracking
13. VT (View Transformation) The views of the target object vary during tracking
14. FM (Fast Motion) The target object moves quickly
3.4 Statistical Analysis
As shown in Table 1, our dataset contains 742 videos, 1,594,474 RGB frames which cover 45 categories of target objects in both indoor and outdoor scenarios. The detailed class names of the target object are shown in Fig (c). 1. We split these videos into training and testing subset, which contains 520 and 222 videos, respectively. For the attribute information, we can find that 549 videos belong to BI (Background Influence) which is the most important challenge in our dataset. OV (Out of View), VT (View Transformation), LI (Low Illumination), and POC (Partial OCclusion) are also involved with more than 400 videos in our dataset.
Figure 2: Representative samples of FELT dataset for long-term frame-event tracking.
4 Methodology
4.1 Overview
As shown in Fig. 3, our long-term frame-event tracking framework, termed AMTTrack, takes the RGB frames and event voxels as input. Then, we extract template voxels/patch and search voxels/patch, and obtain their token representations using the projection layer. These tokens are concatenated and fed into our newly proposed unified associative memory Transformer for feature extraction, interactive learning, and fusion simultaneously. The aggregation of the Hopfield layer and self-attention blocks ensures the associative memory ability for long-term visual tracking. Finally, we obtain the tracking result by using the tracking head which takes the search tokens as the input. More details of our framework will be introduced in the subsequent paragraphs.
4.2 Network Architecture
In this section, we focus on three main modules in our framework, i.e., RGB-Event Encoding Module, Associative Memory Transformer, and Tracking Head.
Input Representation.  Given RGB frames and Event point stream which are denoted as =$\mathcal{I}=\{I_{1},I_{2}\cdots I_{N}\}$ and $\mathcal{E}_{p}=\{e_{1},e_{2}\cdots e_{M}\}$, where $i$ and $j$ denote the index of RGB frame and event point respectively. $e_{j}=\{x,y,t,p\}$ denotes each event point and $(x,y)$ denotes the spatial coordinates. $t$ and $p$ denote timestamps and attributes (polarity) of event point. We first divide the event point stream $\mathcal{E}_{p}$ into the voxel representation $\mathcal{E}_{v}$ in which each voxel grid can be represented as $E_{vg}\in\mathbb{R}^{w\times h\times t}$, where $w$ and $h$ denote the spatial dimension of voxel, and $t$ is the temporal dimension. Then, we extract the template and search image patch and event voxel as the input of our tracker, i.e., $[\mathcal{Z}_{I},\mathcal{Z}_{E},\mathcal{S}_{I},\mathcal{S}_{E}]$. Following CEUTrack [22], we drop the voxel grids whose coordinates are not in the search or template region, and then choose the top 4096 and 1024 voxel grids for search and template region respectively in the sorted voxels (based on the density of each event point) for the event stream. We embed these inputs into their token representations by using projection layers and concatenate them into a unified feature representation $\mathcal{X}=\{\hat{\mathcal{Z}}_{I},\hat{\mathcal{Z}}_{E},\hat{\mathcal{S}}_{I},\hat{\mathcal{S}}_{E}\}$ for tracking. In addition, we also introduce position encoding $\mathcal{P}\in\mathbb{R}^{b\times n\times d}$ and add tokens $\mathcal{X}$ as final features $\mathcal{X}^{\prime}=\mathcal{X}+\mathcal{P}$, where $b$, $n$ and $d$ denote the batch size, number of tokens and embedding dimension respectively.
Figure 3: An overview of our proposed Associative Memory Transformer for long-term frame-event visual tracking. Our tracker (termed AMTTrack) takes the event voxels and RGB images as the input and we divide them into search/template patch/voxels for tracking. After token embedding using projection layers, we concatenate the features and add them with position encoding as the input of our backbone network. It contains Hopfield layer and multi-head self-attention for associative memory storing and retrieval. The enhanced search tokens are fed into the tracking head for target localization.
Associative Memory Transformer.  We designed a unified backbone to mix the RGB-Event tokens which consists of $L$ associative memory Transformer layers. As shown in Fig. 3, the input tokens $\mathcal{X}^{\prime}$ are firstly fed into the Hopfield layer which can learn static queries or stored patterns. It will automatically capture the relations between different training samples and utilize the memorized information to enhance the feature representations for the degenerated modality. More in detail, the output feature $\mathcal{X}^{\prime\prime}\in\mathbb{R}^{b\times n\times d}$ can be obtained via the following operations:
$\mathcal{X}^{\prime\prime}=Softmax\big{(}\beta\mathcal{X}^{\prime}W^{T}_{U}\big{)}W_{C},$ (1)
where $Softmax(\cdot)$ is the Softmax function and $\beta$ is a scaling scalar set to $1/\sqrt{d}$. $W_{U}\in\mathbb{R}^{p\times d}$ and $W_{C}\in\mathbb{R}^{d\times p}$ are two learnable parameters defined in the Hopfield layers, where $p$ is the number of prototypes. Obviously, the Hopfield layer will learn the most representative features and save the stored patterns in $W_{C}$, i.e., each row in matrix $W_{C}$ corresponds to a learned RGB-Event feature. With the associative memory ability, our RGB-Event visual trackers can perform well in the long-term scenarios, as further validated in Section 5.4.
After we obtain $\mathcal{X}^{\prime\prime}$, we feed them into several standard Transformer blocks and each block contains LayerNorm, Multi-Head Self-Attention (MHSA), and Feed Forward layers, as shown in Fig. 3. To be specific, we first transform the $\mathcal{X}^{\prime\prime}$ into $Q=W_{Q}\mathcal{X}^{\prime\prime}$, $K=W_{K}\mathcal{X}^{\prime\prime}$ and $V=W_{V}\mathcal{X}^{\prime\prime}$, where $\{W_{Q},W_{K},W_{V}\}\in\mathbb{R}^{d\times d}$ denote the query, key and value transformation matrices, respectively. Then, we adopt Multi-Head Self-Attention module to capture the long-range relationships of all tokens as:
$\bar{\mathcal{X}}=Softmax(\frac{QK^{T}}{\sqrt{d}})V$ (2)
The output $\bar{\mathcal{X}}$ will be fed into LayerNorm and Feed Forward layers, respectively. Note that, our associative memory Transformer blocks can be stacked for multiple layers, our experiments demonstrate that the better tracking performance can be obtained when $L$ is set as 8.
Tracking Head & Loss Function.  After we obtain the enhanced template and search tokens for both modalities, following OSTrack [30] and CEUTrack [22], we feed the tokens of search patch and voxels into the tracking head for target object localization. The padded sequence of search tokens is firstly re-interpreted into a 2D spatial feature map, then, a fully convolutional network (FCN) which consists of stacked Conv-BN-ReLU layers is employed for each output. More in detail, the tracking head contains three branches that are used for the prediction of classification score, center offset, and the scale of the bounding box, respectively.
For the training of our network, we adopt a combined loss function that contains the focal loss [14] for classification and adopts L1 loss and GIoU loss [20] for bounding box regression, i.e.,
$\mathcal{L}=\lambda_{1}\mathcal{L}_{focal}+\lambda_{2}\mathcal{L}_{1}+\lambda_{3}\mathcal{L}_{GIoU},$ (3)
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are all trade-off parameters. We experimentally set them as 1, 1 and 14, respectively.
5 Experiment
5.1 Datasets and Evaluation Metric
For the evaluation of our tracker, we conduct our experiments on the long-term frame-event tracking dataset FELT. In addition, we also evaluate it on the RGB-Thermal tracking task using LasHeR [16] dataset to demonstrate the generalization of our tracking framework. LasHeR dataset is a large-scale benchmark dataset proposed for RGB-Thermal tracking, which contains 1,224 video sequences (pairs with 730K image pairs). It default divides the dataset into a training and a testing subset which contains 979 and 245 video sequence pairs.
For the evaluation of the trackers, we adopt the widely used Precision Rate (PR), and Success Rate (SR). Specifically speaking, the Precision Rate (PR) illustrates the percentage of frames where the center location error between the object location and ground truth is smaller than a pre-defined threshold (20 pixels is adopted in this work by default). Success Rate (SR) demonstrates the percentage of frames the IoU of the predicted and the ground truth bounding boxes is higher than a given ratio.
5.2 Implementation Details
In the training phase, the template and search images are randomly sampled in a 100-frame window by following CEUTrack [22]. The AdamW [17] optimizer is adopted for the optimization of our network for a total of 50 epochs. The weight decay is 1e-4 and batch size is 32. We set the initial learning rate as 1e-4 and decay a magnitude after the $40^{th}$ epoch. Following CEUTrack, we don’t employ any data augmentation for the training. More details can be found in our source code in the supplementary materials.
Table 3: Tracking results on FELT SOT dataset (SR/PR).
Trackers Source RGB Event RGB+Event FPS
01. MixFormer [4] CVPR22 48.0/61.9 38.9/50.4 48.1/63.0 25
02. AiATrack [10] ECCV22 46.7/58.9 40.3/51.6 46.6/59.5 38
03. STARK [28] ICCV21 45.6/58.2 39.3/50.8 45.7/59.4 42
04. OSTrack [30] ECCV22 45.2/56.3 37.4/46.9 32.5/40.3 105
05. GRM [11] CVPR23 44.7/55.9 39.2/48.9 44.5/55.9 45
06. PrDiMP [6] CVPR20 43.8/54.7 34.9/44.5 43.8/55.2 30
07. DiMP [1] ICCV19 43.3/54.5 37.8/48.5 43.5/55.1 43
08. SuperDiMP [1] - 43.0/53.5 37.8/47.8 43.0/54.2 -
09. TransT [3] CVPR21 42.2/53.0 35.2/45.1 34.6/44.3 50
10. TOMP50 [18] CVPR22 41.2/51.8 37.7/49.2 43.4/55.2 25
11. ATOM [5] CVPR19 37.5/47.0 22.3/28.4 36.2/45.9 30
12. KYS [2] ECCV20 35.9/45.0 22.5/29.5 33.1/42.4 20
13. CEUTrak [22] arXiv23 - - 44.9/56.4 75
14. HRCEUTrack [40] ICCV23 - - 44.8/56.4 75
15. AFNet [32] CVPR23 - - 28.9/36.6 36
AMTTrack (Ours) - - - 45.5/57.2 61
5.3 Comparison on Public Benchmarks
Results on FELT Dataset. As shown in Table 3 and Fig. 4, we report the results of existing SOTA trackers and our trackers on the FELT SOT dataset. To be specific, our baseline tracker CEUTrack [22] achieves $44.9/56.4$ on SR/PR metric, meanwhile, the results can be improved to $45.5/57.2$ when using the Hopfield layers (AMTTrack). This experiment demonstrates that associative memory helps spatial-temporal feature learning for long-term frame-event tracking. These experiments fully validate the effectiveness of our proposed strategies for long-term frame-event single object tracking.
Compared with other state-of-the-art trackers, our proposed AMTTrack ranks the $4^{th}$ place on the FELT benchmark. The first, second, and third-ranked tracker is MixFormer [4], AiATrack [10], and STARK [28] which are also strong trackers developed based on Transformer in the standard visual tracking community. Our proposed algorithm can beat all other trackers in Table 3.
Figure 4: Tracking results on our newly proposed FELT long-term tracking dataset.
Results on LasHeR Dataset. As shown in Table 4, we compare our tracker with some latest state-of-the-art trackers on the large-scale LasHeR dataset, such as CEUTrack [22], ViPT [36], etc. When introducing the Hopfield network layers, our precision rate and success rate can be improved from the original 63.5/49.6 (CEUTrack) to the current 64.8/50.8 (AMTTrack). Obviously, the results on the RGB-T dataset showed a greater improvement compared to the FELT dataset. This experiment demonstrates the effectiveness and generalization of our method on multiple RGB+X based tracking tasks.
Table 4: Tracking results on LasHeR SOT dataset.
Trackers DAPNet ViPT ProTrack TransT OSTrack HMFT SGT++
 [37]  [36]  [29]  [3]  [30]  [35]  [16]
PR 43.1 65.1 53.8 52.4 51.5 43.6 36.5
SR 31.4 52.5 42.0 39.4 41.2 31.3 25.1
Trackers APFNet CAT STARKS50 mfDiMP FANet CEUTrack Ours
 [26]  [15]  [28]  [34]  [38]  [22] -
PR 50.0 45.0 44.9 44.7 44.1 63.5 64.8
SR 36.2 31.4 36.1 34.3 30.9 49.6 50.8
5.4 Ablation Study
Analysis of Hopfield Layers. To further validate the effectiveness of the used Hopfield layer, in this experiment, we set different layers to check its influence. As shown in Table 5, we set 12, 8, and 6 layers and find that slightly better results can be obtained when 8 Hopfield layers are used.
Table 5: Different Hopfield layers on FELT dataset (SR/PR).
Layers 6 8 12
Hopfield Layer 45.4/57.1 45.5/57.2 45.2/56.9
Tracking Results Under Challenging Scenarios. In this paper, we define 14 attributes for the evaluation of visual trackers when facing challenging factors. As shown in Fig. 5, current SOTA trackers achieve a relatively high score when meeting attributes like SV (Scale Variation), NMO (No Motion), DEF (Deformation), VT (View Transformation) and IV (Illumination Variation). However, the performance is still poor when tracking the ST (Small Targets), FM (tracking in Fast Motion) and OE (Over Exposure). Our proposed FELT dataset provides a better platform to boost the related research on frame-event long-term single object tracking problem.
Figure 5: Tracking results (SR) under 12 challenging factors.
Analysis on Different Input Modalities. As shown in Table 3, we re-train SOTA trackers using three different sets, i.e., RGB frames, Event streams, and RGB+Event data. We can find that current SOTA trackers perform better on RGB frames than event streams. For example, MixFormer achieves $48.0/61.9$ based on RGB frames, however, only obtains $38.9/50.4$ on Event data. These comparisons demonstrate that tracking using event data is still a challenging task. When combining the two modalities for tracking, the overall performance can be improved for some trackers such as MixFormer, Stark, PrDiMP, etc. However, different from regular multi-modal tracking tasks, such as RGB-Thermal, the results of some trackers are decreased when using both RGB and Event data, such as OSTrack, ATOM, KYS, etc. These results demonstrate that the intelligent fusion strategy for RGB and Event is still necessary and urgent.
Figure 6: Tracking results of our AMTTrack and other trackers.
Efficiency Analysis. Our proposed tracker achieves 61 FPS when tracking on the FELT dataset. The model size of our tracker is 542.27 MB and it contains 117.79M parameters.
Figure 7: Activation maps obtained using our AMTTrack.
5.5 Visualization
In addition to the above quantitative analysis, we also present some visual demonstrations to better help readers understand our tracking algorithm. As shown in Fig. 6, we can find that our proposed tracker can achieve robust tracking results even compared with state-of-the-art visual trackers. Also, our response maps can always focus on the true target object, as shown in Fig. 7. Therefore, we can draw the conclusion that our proposed associative memory Transformer works well for the frame-event long-term visual tracking task.
5.6 Limitation Analysis
Our proposed tracker achieves good performance on FELT and other tracking datasets, however, its overall performance is still limited compared with other SOTA trackers. Our tracker adopts a unified backbone for RGB-Event feature extraction and fusion, therefore, we attempt to improve the backbone by pre-training on unlabelled RGB-Event data.
6 Conclusion
In this work, we propose a new long-term and large-scale frame-event single object tracking dataset, termed FELT. It contains 742 videos and 1,594,474 RGB frames and event stream pairs and is also the largest frame-event tracking dataset to date. To build a comprehensive benchmark, we also re-train and evaluate 15 baseline trackers on our newly proposed FELT SOT dataset, which will be a better platform for future works to compare. In addition, we find that the RGB frames and event streams are naturally incomplete data due to the influence of challenging factors and spatially sparse event data. To address this issue, in this work, we propose a novel associative memory Transformer network by introducing modern Hopfield layers into multi-head self-attention blocks to fuse both RGB and event data. Extensive experiments on multiple event-based tracking datasets fully validated the effectiveness of our model. In our future works, we will consider further improving the overall long-term tracking results by exploiting the dense temporal resolution of event data via asynchronous fusion modules.